# A Gaussian integral formula for the Hermite polynomials: Combinatorics, Asymptotics and Applications

### v1

@@ 1 Introduction to the formula and its uses
To the best of our knowledge, this formula was first written about in [Wit00]. For probabilists, who are adept at manipulating expected values with Gaussians, this formula can be incredibly useful. For example, the identity for the value at 0,
$H\!\!\operatorname{e}_{n}(0)=\mathbb{E}[(0+\imath Z)^{n}]=\begin{cases}(-1)^{n/2}(n-1)!!&\text{ for }n\text{ even}\\ 0&\text{ for }n\text{ odd}\end{cases},$ (3)
is an immediate consequence. (See also Section 2 for a combinatorial interpretation of $\mathbb{E}[Z^{n}]=(n-1)!!$)
@@
Note also that by plugging in $x=0$ into (4) and using (3), one obtains an explicit sum formula for $H\!\!\operatorname{e}_{n}$. (See again Section 2 for a combinatorial interpretation.)
@@ 1.1 Roadmap of this article
The ethos of this article is to use the Gaussian integral formula (2) of Definition 1.1 as our starting definition for the Hermite polynomials, and then to showcase alternate derivations of many of the properties of Hermite polynomials and derive several applications. Most importantly, we also show throughout how this formula allows one to easily generalize the Hermite polynomials and their properties to multiple dimensions where possible. Below is a brief summary of each section.
@@
  3. §4 $\Psi_{n}(x)$ and $K_{n}(x,x)$
    We define the oscillator wave function $\Psi_{n}(x)$ and the Hermite kernel $K_{n}(x,x)$, which are a common way the Hermite polynomials appear in applications. We prove several properties of these using the results of Section 3, including convergence (after appropriate rescaling) to the arcsine law (see Figure 7), semi-circle law (see Figure 8) and the Airy limit.
  4. §5 Random matrices
@@
Each section can be understood without having understood any other section, and sections only rarely mention results from other sections.
However, there is a logical progression within $\S$ 3 $\to\S$ 4 $\to\S$ 5 , where each subsequent section uses a pre-packaged result from the previous section. In $\S$ 3 the actual detailed asymptotic analysis of $H\!\!\operatorname{e}_{n}$ is calculated “by hand”, then in $\S$ 4, the asymptotics are repackaged to give nice results for $\Psi_{n}(x)$ and $K_{n}(x,x)$, and in $\S$ 5, those results for convergence of functions are used to prove limit theorems for random matrices and Dyson Brownian motion.
@@ 1.2 Generalization to arbitrary variance $\widetilde{H}_{n}(x,t)$ and its dual $H_{n}(x,t)$
$\widetilde{H}_{n}(x,t)=t^{n/2}\widetilde{H}_{n}\left(\frac{x}{\sqrt{t}},1\right)=t^{n/2}H\!\!\operatorname{e}\left(\frac{x}{\sqrt{t}}\right),\quad H_{n}(x,t)=t^{-n/2}H_{n}\left(\frac{x}{\sqrt{t}},1\right)=t^{-n/2}H\!\!\operatorname{e}_{n}\left(\frac{x}{\sqrt{t}}\right)$ (6)
@@
We can also easily generalize to multi-variable Hermite polynomials using this formula, as suggested in [Wit00]. The generalization to $d$ dimensions is to simply replace the single Gaussian $Z\sim\mathcal{N}(0,1)$ with a multi-variate Gaussian $\vec{Z}\sim\mathcal{N}(\vec{0},V)$ for some $d\times d$ covariance matrix $V$. The index of the Hermite polynomial is now a $d$-tuple $(n_{1},\ldots,n_{d})$, and the Hermite polynomial is the the expectation.
@@ 1.4 A warning regarding the Physicist’s Hermite polynomials $H_{n}(x)$
WARNING: Do not confuse the notation for the multivariable (probabilists) Hermite polynomials $H_{\vec{n}}(\vec{x};V)$ from Definition 1.5 with the single variable physicists Hermite polynomials $H_{n}(x)$. Everything in this article is written using only the probabilist’s Hermite polynomials $H\!\!\operatorname{e}_{n}(x)$, $\widetilde{H}_{\vec{n}}(\vec{x};V)$ or $H_{\vec{n}}(\vec{x};V)$. We do not mention the physicists Hermite polynomials $H_{n}$ in this article except to warn the reader that they are different!
@@ 2 Combinatorics of the Hermite polynomials
Note that previous authors have derived many combinatorial interpretations of the Hermite polynomials, see [Dra09, LY89] and references therein. However, the starting point in those presentation is the recurrence relation for the Hermite polynomials, rather than the Gaussian integral formula. Our presentation here gives an extra level of meaning to the combinatorics. For example, in those papers, the authors declare that certain edge weight should be 1-1, and prove that this recovers the Hermite polynomials by induction with the recurrence relation. In contrast, our starting point realizes the 1-1 as arising from an expectation of imaginary Gaussians squared: $1=\mathbb{E}[(\imath Z)(\imath Z)]$. This gives a new probabilistic meaning to the combinatorial methods. We also generalize the proofs to the multivariate Hermite polynomials.
@@ 2.1 Tool: Wick formula/Isserlis’ theorem for Gaussian expectations
The entry point into combinatorics is the Wick formula/Isserlis’ theorem, which states that a Gaussian expectation for mean zero Gaussians can be seen as a sum over pairings. Namely, if $Z_{1},\ldots,Z_{n}$ are a collection of mean zero, jointly Gaussian random variables, then the expectation of the product $Z_{1}\cdots Z_{n}$ can be written in terms of the individual correlations $\mathbb{E}[Z_{a}Z_{b}]$ as:
@@ 2.3 Combinatorics of the multivariable Hermite polynomials $\widetilde{H}_{\vec{n}}(\vec{x},V)$
$\widetilde{H}_{3,1}(x_{1},x_{2};V)=x_{1}^{3}x_{2}-3\varrho x_{1}^{2}-3x_{1}x_{2}+3\varrho$
@@
$\widetilde{H}_{2,2}(x_{1},x_{2};V)=x_{1}^{2}x_{2}^{2}-x_{1}^{2}-x_{2}^{2}-4\varrho x_{1}x_{2}+2+\varrho^{2}$
@@ 3 Asymptotics for the Hermite polynomials as $n\to\infty$
The Gaussian integral formula (7) also yields an easy entry point to the large $n$ limit of the Hermite polynomials because it is naturally set up for the Laplace approximation method for integrals. Other existing approaches to the Hermite asymptotics instead use the derivative formula and then must go through the Cauchy integral formula and residue analysis to the turn the derivative into a complex integral e.g. Section 3 of [AGZ09] and [Sze75, Section 8] . For this reason, unlike the typical statement, the asymptotic formulas we derive here does not involve the factorial $n!$ (which appears due to the Cauchy integral formula), thereby giving “clean” asymptotics without the need for Stirling’s formula.
@@
Figure 4: The approximation formulas for various regimes from Propositions 3.5, 3.1 and 3.6 compared to $H\!\!\operatorname{e}_{n}(x)$ in the case $n=25$. The $y$-axis is a modified log scale, so that the sign of each $y$ value is preserved and 0 is mapped to 0; the map is $y\to sgn(y)\log_{10}(|y|+1)$. The relatively simple asymptotics when $x=O(1)$ and $x=O(n)$ are only accurate when $x\ll\sqrt{n}$ and $x\gg\sqrt{n}$ (and in fact can also be recovered as limits of the more complicated $x=O(\sqrt{n})$ asymptotics). The $x=O(\sqrt{n})$ asymptotics are divided into three regimes; $x=2\sqrt{n}\cos(\alpha)<2\sqrt{n}$ for small $x$ values, $x=2\sqrt{n}\cosh(\gamma)>2\sqrt{n}$ for large $x$ values, and $x=2\sqrt{n}+n^{-1/6}u$ for $x$ values on the edge near $x=2\sqrt{n}$. Notice from the zoomed insert that the small-x and large-x approximations are both singular at the edge $x=2\sqrt{n}$, and this is the regime where the edge approximation with the Airy function is accurate.
@@ 3.1 Approximation to $H\!\!\operatorname{e}_{n}(x)$ when $x=O(n)$
Now observe the limit $\lim_{n\to\infty}(1+i\frac{Z}{cn})^{n}=e^{iZ/c}$ almost surely in $Z$ (see e.g. Theorem 3.4.2 in [Dur10]), and in light of the expectation $\mathbb{E}[e^{\beta Z}]=e^{\frac{1}{2}\beta^{2}}$ we find that $\mathbb{E}[(1+i\frac{Z}{cn})^{n}]=e^{-\frac{1}{2c^{2}}}(1+o(1))$, whence the result follows.
@@ 3.2 Tool: Laplace approximation for integrals
[Laplace Approximation] Suppose $g(x)$ is a continuous function, and suppose $f(x)$ is a function with a unique global maximum at $x^{\ast}$ with zero-derivative there, $f^{\prime}(x^{\ast})=0$, and strictly negative second derivative there $f^{\prime\prime}(x^{\ast})<0$. Suppose also $x^{\ast}\in(a,b)$ is some interval containing $x^{\ast}$ as an interior point. Then, under certain regularity conditions on $f$ (see e.g. Theorem 3.5.3 from [AGZ09] for a precise statement), we have the following Laplace approximation of the integral as $n\to\infty$,
$\intop_{a}^{b}e^{nf(x)}g(x)\mathrm{d}x=e^{nf(x^{\ast})}g(x^{\ast})\sqrt{\frac{2\pi}{n|f^{\prime\prime}(x^{\ast})|}}(1+o(1)).$
@@
The proof idea of the Laplace approximation can be more powerful than the mere statement of the result, so we briefly elucidate it here. The idea is to approximate $f$ around its global maximum by Taylor’s theorem $f(x)=f(x^{\ast})-\frac{1}{2}|f^{\prime\prime}(x^{\ast})|(x-x^{\ast})^{2}+\epsilon_{2}\left(x-x^{\ast}\right)$, and then do a change of variables $\frac{s}{\sqrt{n}}=x-x^{\ast}$, so that
$\intop_{x^{\ast}-c/\sqrt{n}}^{x^{\ast}+c/\sqrt{n}}e^{nf(x)}\mathrm{d}x=e^{nf(x^{\ast})}\intop_{-c}^{c}e^{-\frac{1}{2}|f^{\prime\prime}(x^{\ast})|s^{2}+n\epsilon_{2}\left(\frac{s}{\sqrt{n}}\right)}\frac{\mathrm{d}s}{\sqrt{n}}.$
Then, by using the technical regularity conditions of the function $f$, one argues that Taylor error term $n\epsilon_{2}(\frac{s}{\sqrt{n}})$ goes to zero fast enough to be negligeable (For example, if $f$ has bounded third derivative, one has $\epsilon_{2}(y)=O(y^{3})$ and $n\epsilon_{2}(\frac{s}{\sqrt{n}})$ goes to zero like $\frac{1}{\sqrt{n}}$ ) , whence what remains is the Gaussian integral $\intop e^{-\frac{1}{2}|f^{\prime\prime}(x^{\ast})|s^{2}}\mathrm{d}s=\sqrt{\frac{2\pi}{|f^{\prime\prime}(x^{\ast})|}}$. To complete the argument, one also needs to argue that the integral outside of this neighbourhood of $x^{\ast}$ is exponentially smaller than $e^{nf(x^{\ast})}$ and use continuity of $g$ to bring $g(x)$ along for the ride. (see e.g. Theorem 3.5.3 from [AGZ09] where this is carried out)
This idea illustrates a few important concepts:

### v2

@@ Abstract
4. An analysis of a phase transition in the spiked GUE random matrix as the top eigenvalue goes from well-separated to attached to the bulk, analogous to the BBP phase transition.
@@ 1 Introduction to the formula and its uses
This formula has appeared in many places, for example Exercise 6.8 in [Temme] or (4.9) in [janson]. Integral formulas for other orthogonal polynomials also exist, such as (4.5.17) in [ismael].
For probabilists, who are adept at manipulating expected values with Gaussians, this formula can be exceptionally useful. For example, the identity for the value of $H\!\!\operatorname{e}_{n}$ at $x=0$ follows from $\mathbb{E}[Z^{n}]=(n-1)!!$ when $n$ is even,
$H\!\!\operatorname{e}_{n}(0)=\mathbb{E}[(0+\imath Z)^{n}]=\begin{cases}(-1)^{n/2}(n-1)!!&\text{ for }n\text{ even}\\ 0&\text{ for }n\text{ odd}\end{cases}.$ (3)
See also Section 2 for a combinatorial interpretation of $\mathbb{E}[Z^{n}]=(n-1)!!$.
@@
One can also generalize this approach to more summands to get expressions for $H\!\!\operatorname{e}_{n}(x_{1}+\ldots+x_{m})$, see Theorem 5.5. of [MollVignat]. Note also that by plugging in $x=0$ into (4) and using (3), one obtains an explicit sum formula for $H\!\!\operatorname{e}_{n}$. (See again Section 2 for a combinatorial interpretation.)
@@ 1.1 Roadmap of this article
The ethos of this article is to use the Gaussian integral formula (2) of Definition 1.1 as our starting definition for the Hermite polynomials, and then to showcase alternate derivations of many of the properties of Hermite polynomials and obtain several applications. We also show throughout how this formula allows one to easily generalize the Hermite polynomials and their properties to multiple dimensions where possible. Below is a brief summary of each section.
@@
  3. §4 $\Psi_{n}(x)$ and $K_{n}(x,y)$
    We define the oscillator wave function $\Psi_{n}(x)$ and the Hermite kernel $K_{n}(x,y)$, which are a common way the Hermite polynomials appear in applications. We prove several properties of these using the results of Section 3, including convergence (after appropriate rescaling) to the arcsine law (see Figure 8), semi-circle law (see Figure 9) and the Airy limit.
  4. §5 Random matrices I
@@
  5. §6 Random matrices II
    It turns out that a rank one perturbation to the GUE (sometimes called a “spiked random matrix”) also has explicit formulas given purely in terms of Hermite polynomials. By analyzing these in detail, we find a phase transition where the top eigenvalue is well separated from the bulk and then another phase where it is glued to the bulk. This is analogous to the so-called “BBP phase transition” [BBP] that appears in the sample covariance matrix.
@@
Each section can be understood without having understood the detailed proof of the other section. However, there is a logical progression within $\S$3 $\to\S$4 $\to\S$5 $\to\S$6, where each subsequent section uses pre-packaged results from previous sections. In $\S$3 the actual detailed asymptotic analysis of $H\!\!\operatorname{e}_{n}$ is calculated “by hand”, then in $\S$4, the asymptotics are repackaged to give nice results for $\Psi_{n}(x)$ and $K_{n}(x,x)$. In $\S$5, those results for convergence of functions are used to prove limit theorems for random matrices and Dyson Brownian motion and extended to spiked Dyson Brownian motion in §6.
@@ 1.2 Generalization to arbitrary variance $\widetilde{H}_{n}(x,t)$ and its dual $H_{n}(x,t)$
$\widetilde{H}_{n}(x,t)=t^{n/2}\widetilde{H}_{n}\left(\frac{x}{\sqrt{t}},1\right)=t^{n/2}H\!\!\operatorname{e}_{n}\left(\frac{x}{\sqrt{t}}\right),\quad H_{n}(x,t)=t^{-n/2}H_{n}\left(\frac{x}{\sqrt{t}},1\right)=t^{-n/2}H\!\!\operatorname{e}_{n}\left(\frac{x}{\sqrt{t}}\right)$ (6)
@@
Hermite polynomials naturally generalize to higher dimensions, see for example [Appell, AKF]. As suggested in [Withers2000], the Gaussian expectation formula (2) can also be used here. The generalization to $d$ dimensions is to simply replace the single Gaussian $Z\sim\mathcal{N}(0,1)$ with a multi-variate Gaussian $\vec{Z}\sim\mathcal{N}(\vec{0},V)$ for some $d\times d$ covariance matrix $V$. The index of the Hermite polynomial is now a $d$-tuple $(n_{1},\ldots,n_{d})$, and the Hermite polynomial is the the expectation.
@@ 1.4 A warning regarding the Physicist’s Hermite polynomials $H_{n}(x)$
WARNING: Do not confuse the notation for the multivariable (probabilists) Hermite polynomials $H_{\vec{n}}(\vec{x};V)$ from Definition 1.5 with the single variable physicists Hermite polynomials $H_{n}(x)$. Everything in this article is written using only the probabilist’s Hermite polynomials $H\!\!\operatorname{e}_{n}(x)$, $\widetilde{H}_{\vec{n}}(\vec{x};V)$ or $H_{\vec{n}}(\vec{x};V)$. We do not mention the physicists Hermite polynomials $H_{n}$ in this article except in this section in order to warn the reader that they are different!
@@ 2 Combinatorics of the Hermite polynomials
Note that previous authors have derived many combinatorial interpretations of the Hermite polynomials, see [HermiteCombinatorics, PolynomialCombinatorics, janson] and references therein. However, the starting point in those presentation is the recurrence relation for the Hermite polynomials, rather than the Gaussian integral formula. Our presentation here gives an extra level of meaning to the combinatorics. For example, in those papers, the authors declare that certain edge weight should be 1-1, and prove that this recovers the Hermite polynomials by induction with the recurrence relation. In contrast, our starting point realizes the 1-1 as arising from an expectation of imaginary Gaussians squared: $1=\mathbb{E}[(\imath Z)(\imath Z)]$. This gives a new probabilistic meaning to the combinatorial methods. We also generalize the proofs to the multivariate Hermite polynomials.
@@ 2.1 Tool: Wick formula/Isserlis’ theorem for Gaussian expectations
The entry point into combinatorics is the Wick formula/Isserlis’ theorem, which states that a Gaussian expectation for mean zero Gaussians can be seen as a sum over pairings. (These also sometimes go by the name of “Feynman Diagrams” see e.g. [janson] Sections 1.5 and 3.1. Be warned that “Feynman Diagrams” can also refer to an entirely different set of diagrams from subatomic physics.)
Consider as follows: if $Z_{1},\ldots,Z_{n}$ are a collection of mean zero, jointly Gaussian random variables, then the expectation of the product $Z_{1}\cdots Z_{n}$ can be written in terms of the individual correlations $\mathbb{E}[Z_{a}Z_{b}]$ as:
@@ 2.3 Combinatorics of the multivariable Hermite polynomials $\widetilde{H}_{\vec{n}}(\vec{x},V)$
Remark 2.1.
These combinatorial expansions are intimately related to the combinatorial expansions for Wick products, see e.g. [janson] Section 3.1.
@@
$\widetilde{H}_{3,1}(x_{1},x_{2};V)=\mathbb{E}[(x_{1}+\imath Z_{1})^{3}(x_{2}+\imath Z_{2})]=x_{1}^{3}x_{2}-3\varrho x_{1}^{2}-3x_{1}x_{2}+3\varrho$
@@
$\widetilde{H}_{2,2}(x_{1},x_{2};V)=\mathbb{E}[(x_{1}+\imath Z_{1})^{2}(x_{2}+\imath Z_{2})^{2}]=x_{1}^{2}x_{2}^{2}-x_{1}^{2}-x_{2}^{2}-4\varrho x_{1}x_{2}+2+\varrho^{2}$
@@ 3 Asymptotics for the Hermite polynomialsas $n\to\infty$
The Gaussian integral formula (7) also yields an easy entry point to the large $n$ limit of the Hermite polynomials because it is naturally set up for the Laplace approximation method for integrals. Other existing approaches to the Hermite asymptotics instead use the derivative formula and then must go through the Cauchy integral formula and residue analysis to the turn the derivative into a complex integral e.g. Section 3 of [AGZ] and [Szego1975, Section 8] . For this reason, unlike the typical statement, the asymptotic formulas we derive here does not involve the factorial $n!$ (which appears due to the Cauchy integral formula), thereby giving “clean” asymptotics without the need for Stirling’s formula.
@@
Figure 4: The approximation formulas for various regimes from Propositions 3.6, 3.1 and 3.7 compared to $H\!\!\operatorname{e}_{n}(x)$ in the case $n=25$. The $y$-axis is a modified log scale, so that the sign of each $y$ value is preserved and 0 is mapped to 0; the map is $y\to sgn(y)\log_{10}(|y|+1)$. The relatively simple asymptotics when $x=O(1)$ and $x=O(n)$ are only accurate when $x\ll\sqrt{n}$ and $x\gg\sqrt{n}$ (and in fact can also be recovered as limits of the more complicated $x=O(\sqrt{n})$ asymptotics). The $x=O(\sqrt{n})$ asymptotics are divided into three regimes; $x=2\sqrt{n}\cos(\alpha)<2\sqrt{n}$ for small $x$ values, $x=2\sqrt{n}\cosh(\gamma)>2\sqrt{n}$ for large $x$ values, and $x=2\sqrt{n}+n^{-1/6}u$ for $x$ values on the edge near $x=2\sqrt{n}$. Notice from the zoomed insert that the small-x and large-x approximations are both singular at the edge $x=2\sqrt{n}$, and this is the regime where the edge approximation with the Airy function is accurate.
@@ 3.1 Approximation to $H\!\!\operatorname{e}_{n}(x)$ when $x=O(n)$
Now observe the limit $\lim_{n\to\infty}(1+i\frac{Z}{cn})^{n}=e^{iZ/c}$ almost surely in $Z$ (see e.g. Theorem 3.4.2 in [durrettPTE]), and in light of the expectation $\mathbb{E}[e^{\beta Z}]=e^{\frac{1}{2}\beta^{2}}$ we find that $\mathbb{E}[(1+i\frac{Z}{cn})^{n}]=e^{-\frac{1}{2c^{2}}}(1+o(1))$, whence the result follows.
@@ 3.2 Tool: Laplace approximation for integrals
[Laplace Approximation] Suppose $g(x)$ is a continuous function, and suppose $f(x)$ is a function with a unique global maximum at $x^{\ast}$ with zero-derivative there, $f^{\prime}(x^{\ast})=0$, and strictly negative second derivative there $f^{\prime\prime}(x^{\ast})<0$. Suppose also $x^{\ast}\in(a,b)$ is some interval containing $x^{\ast}$ as an interior point. Then, under certain regularity conditions on $f$ (see e.g. Theorem 3.5.3 from [AGZ] for a precise statement), we have the following Laplace approximation of the integral as $n\to\infty$,
$\intop_{a}^{b}e^{nf(x)}g(x)\mathrm{d}x=e^{nf(x^{\ast})}g(x^{\ast})\sqrt{\frac{2\pi}{n|f^{\prime\prime}(x^{\ast})|}}(1+o(1)).$ (41)
@@
The proof idea of the Laplace approximation can be more powerful than the mere statement of the result, so we briefly elucidate it here. The idea is to approximate $f$ around its global maximum by Taylor’s theorem $f(x)=f(x^{\ast})-\frac{1}{2}|f^{\prime\prime}(x^{\ast})|(x-x^{\ast})^{2}+\epsilon_{2}\left(x-x^{\ast}\right)$, and then do a change of variables $\frac{s}{\sqrt{n}}=x-x^{\ast}$, so that for any constants $c_{a}<0<c_{b}$ one has
$\intop_{x^{\ast}+c_{a}/\sqrt{n}}^{x^{\ast}+c_{b}/\sqrt{n}}e^{nf(x)}\mathrm{d}x=e^{nf(x^{\ast})}\intop_{c_{a}}^{c_{b}}e^{-\frac{1}{2}|f^{\prime\prime}(x^{\ast})|s^{2}+n\epsilon_{2}\left(\frac{s}{\sqrt{n}}\right)}\frac{\mathrm{d}s}{\sqrt{n}}.$ (42)
Then, by using the technical regularity conditions of the function $f$, one argues that Taylor error term $n\epsilon_{2}(\frac{s}{\sqrt{n}})$ goes to zero fast enough to be negligible (For example, if $f$ has bounded third derivative, one has $\epsilon_{2}(y)=O(y^{3})$ and $n\epsilon_{2}(\frac{s}{\sqrt{n}})$ goes to zero like $\frac{1}{\sqrt{n}}$ ) , whence one remains with the Gaussian integral $\intop_{c_{a}}^{c_{b}}e^{-\frac{1}{2}|f^{\prime\prime}(x^{\ast})|s^{2}}\mathrm{d}s$. As $c_{a}\to$-$\infty$ and $c_{b}\to+\infty$ , this converges to $\sqrt{\frac{2\pi}{|f^{\prime\prime}(x^{\ast})|}}$ which is the term that appears on the RHS of (41). To complete the argument, one also needs to argue that the integral outside of this neighbourhood of $x^{\ast}$ is exponentially smaller than $e^{nf(x^{\ast})}$ and use continuity of $g$ to bring $g(x)$ along for the ride. (see e.g. Theorem 3.5.3 from [AGZ] where these details are carried out and a rigourous proof is given.)
This proof sketch illustrates a few important concepts:
@@
    Since the dominant contribution happens near the maximizer $x^{\ast}$, one can also give limits that give the asymptotic behaviour on windows that zoom in around $x^{\ast}$. For example, by changing variables $z=\sqrt{|f^{\prime\prime}(x^{\ast})|}s$ in the integral of (42), one sees that for $c_{a}<0<c_{b}$ one has the limit
    $\intop_{x^{\ast}+c_{a}/\sqrt{n}}^{x^{\ast}+c_{b}/\sqrt{n}}e^{nf(x)}g(x)\mathrm{d}x=e^{nf(x^{\ast})}g(x^{\ast})\sqrt{\frac{2\pi}{n|f^{\prime\prime}(x^{\ast})|}}F_{\Phi}\Bigg(c_{a}\sqrt{|f^{\prime\prime}(x^{\ast})|},c_{b}\sqrt{|f^{\prime\prime}(x^{\ast})|}\Bigg)(1+o(1)),$ (43)
    where $F_{\Phi}(u,v)=\frac{1}{\sqrt{2\pi}}\intop_{u}^{v}e^{-\frac{1}{2}z^{2}}\mathrm{d}z=\mathbb{P}_{Z\sim\mathcal{N}(0,1)}(u<Z<v)$ is the cumulative probability function for the standard Gaussian random variable.
  4. 4.
@@
Remark 3.5.
While we do not prove this here, the formula given for the limiting behaviour of $n!$ is in fact a lower bound. It will be helpful for us to use the following upper and lower bounds that follow from [Robbins-Stirling]: for any $n\in\mathbb{N}$,
$\sqrt{2\pi n}\left(\frac{n}{e}\right)^{n}<n!\leq e\sqrt{n}\left(\frac{n}{e}\right)^{n}.$